OpenStudy (anonymous):

How do I solve: 2^(4x) = 9^(x-1)?

OpenStudy (anonymous):

ln

OpenStudy (anonymous):

\[4x \ln 2=(x-1)\ln 9\] \[\frac{ \ln 2 }{ \ln 9 }=\frac{ x-1 }{ 4x }\]

OpenStudy (anonymous):

this is what I did so far: 4x(x-1) = log9/log2 4x^2-4x = log9/log2 4x^2-4x-(log9/log2)=0?

OpenStudy (anonymous):

and then what happens? o.O

OpenStudy (anonymous):

\[\frac{ln2}{ln9}4x-x=-1\] \[x(\frac{ln2}{ln9}4-1)=-1\] \[x=\frac{-1}{(\frac{ln2}{ln9}4-1)}\]

OpenStudy (anonymous):

basically throw all the xs to one side, factor the x, and then divide the other side by what did not get factored by x.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ohh, i see. alright, thank you!

OpenStudy (anonymous):

ALTERNATIVELY\[(2^4)^x=9^x9^{-1}\] \[(\frac{ 2^4 }{ 9 })^x=\frac{ 1 }{ 9 }\] \[x \rightarrow \log_{\frac{ 16 }{ 9 }}\frac{ 1 }{ 9 }\]